# Survey of Computerized Adaptive Testing: A Machine Learning Perspective

### v1

@@ 8.1. Datasets
To rigorously assess the efficacy of a CAT solution, it is essential to deploy diverse datasets that can effectively challenge and evaluate the system’s capabilities. Three types of data can be employed for this evaluation:
(1) Educational Assessment Datasets: These datasets originate from actual standardized tests such as the SAT, GRE, or GMAT. They encompass a broad spectrum of questions covering multiple subjects, coupled with responses from a substantial pool of test-takers. Such datasets are invaluable for gauging the CAT system’s capacity. However, due to stringent privacy laws, intellectual property rights, and the protective measures taken by educational testing companies regarding their questions and datasets, most of these are proprietary and not easily accessible.
(2) Real-World Educational Data: This category includes data collected from educational environments in practice, such as schools, universities, and online learning platforms. It provides insights into how examinees interact with educational content and assessments in a natural setting. The data may encompass examinee information, performance responses, learning behaviors, question characteristics, etc. We have open-sourced a comprehensive education-related dataset library: https://github.com/bigdata-ustc/EduData. It includes a range of publicly available datasets along with previously private datasets. Additionally, we have provided a detailed data analysis to support further CAT research and application in educational settings. The following are introductions to several commonly used datasets, and more datasets can be found at the EduData gitHub link.
@@ 9. Opportunities for Future Research
Future research should harness machine learning to enhance the multi-dimensionality of the assessment process. This involves not only the traditional response patterns but also the nuanced analysis of process data such as response times and mouse movements, which can provide insights into an examinee’s problem-solving strategies and levels of engagement. Moreover, integrating learning data, such as an examinee’s prior interactions with educational materials, can offer a longitudinal perspective on their learning trajectory and readiness for new concepts. Additionally, the analysis of educational content, encompassing textual, visual, and auditory materials, allows for a richer understanding of how examinees interact with multifaceted information. Such machine learning-driven approaches promise to refine CAT systems comprehensively, enabling them to deliver assessments that are not just accurate reflections of an examinee’s proficiency but also predictive of their potential for future learning.

### v2

@@ 8.1. Datasets
To rigorously assess the efficacy of a CAT solution, it is essential to deploy diverse datasets that can effectively challenge and evaluate the system’s capabilities. Three types of data can be employed for this evaluation: (1) Educational Assessment Datasets: These datasets originate from actual standardized tests such as the SAT, GRE, or GMAT. They encompass a broad spectrum of questions covering multiple subjects, coupled with responses from a substantial pool of test-takers. Such datasets are invaluable for gauging the CAT system’s capacity. However, due to stringent privacy laws, intellectual property rights, and the protective measures taken by educational testing companies regarding their questions and datasets, most of these are proprietary and not easily accessible. (2) Real-World Educational Data: This category includes data collected from educational environments in practice, such as schools, universities, and online learning platforms. It provides insights into how examinees interact with educational content and assessments in a natural setting. The data may encompass examinee information, performance responses, learning behaviors, question characteristics, etc. We have open-sourced a comprehensive education-related dataset library: https://github.com/bigdata-ustc/EduData. It includes a range of publicly available datasets along with previously private datasets. Additionally, we have provided a detailed data analysis to support further CAT research and application in educational settings. The following are introductions to several commonly used datasets, and more datasets can be found at the EduData gitHub link.
@@ 9. Opportunities for Future Research
Future research should harness machine learning to enhance the multi-dimensionality of the assessment. This involves not only the traditional response patterns but also the nuanced analysis of process data such as response times and mouse movements, which can provide insights into an examinee’s problem-solving strategies and levels of engagement. Moreover, integrating learning data, such as the examinee’s prior interactions with materials, can offer a longitudinal perspective on their learning trajectory and readiness for new concepts. Additionally, the analysis of educational content, encompassing textual, visual, and auditory materials (Xiao et al., 2022; Shi et al., 2014), allows for a richer understanding of how examinees interact with multifaceted information. Such machine learning-driven approaches promise to refine CAT systems comprehensively, enabling them to deliver assessments that are not just accurate reflections of an examinee’s proficiency but also predictive of their potential for future learning.